OpenStudy (anonymous):

21. What is the length of the major axis? 22. What is the length of the minor axis? 23. What is the sum of the focal radii? 24. What are the coordinates of the foci?

OpenStudy (anonymous):

\[\frac{ (x-(-4))^{2} }{ 12 }\frac{ (y-1)^{2} }{ 8 } =1\]

OpenStudy (jdoe0001):

any confusing parts?

OpenStudy (anonymous):

isnt the circle in the wrong place? i can't see how far it is from the center

OpenStudy (jdoe0001):

my picture, yes, it was just a representation of a horizontal ellipse

OpenStudy (anonymous):

i just can't see how far it is from the center. i know what you're talking about.. but i don't know the focal points so i have no clue how far it is..

OpenStudy (jdoe0001):

the distance from the center to either focus is "c", \(\bf c=\sqrt{a^2-b^2}\) its sum would be c + c

OpenStudy (anonymous):

oh ok..

OpenStudy (jdoe0001):

\(\bf a=\sqrt{12}\qquad b=\sqrt{8}\\ \quad \\ a^2= 12\qquad b^2=8\)

OpenStudy (anonymous):

thanks ..

OpenStudy (anonymous):

:)

OpenStudy (jdoe0001):

\(\bf a^2=\sqrt{12}\cdot \sqrt{12}\implies \sqrt{12^2}\implies 12\\ \quad \\ b^2=\sqrt{8}\cdot \sqrt{8}\implies \sqrt{8^2}\implies 8\)

OpenStudy (jdoe0001):

the "a" and "b" components are the root version of the denominators

OpenStudy (anonymous):

so 4 / 4

OpenStudy (anonymous):

1

OpenStudy (jdoe0001):

for... ... what part again?

OpenStudy (anonymous):

c / c

OpenStudy (anonymous):

\[c \div c\]

OpenStudy (anonymous):

honestly i don't know

OpenStudy (jdoe0001):

\(\bf c=\sqrt{a^2-b^2}\implies c=\sqrt{12-8}\implies c=\sqrt{4}\implies c=2\)

OpenStudy (anonymous):

thats the major?

OpenStudy (jdoe0001):

... no... that's the "c" component the major axis is "a" units from the center of the ellipse, thus a + a is how long it's the minor is "b" units from the center, thus b + b is how long it's

OpenStudy (jdoe0001):

notice your "a" and "b" \(\bf \bf a=\sqrt{12}\qquad b=\sqrt{8}\\ \quad \\ a^2= 12\qquad b^2=8\)

OpenStudy (anonymous):

omg.. im an idiot... what's c..

OpenStudy (jdoe0001):

"c" is the distance form the vertex to the focii

OpenStudy (jdoe0001):

\(\bf a+a\implies \sqrt{12}+\sqrt{12}\qquad \qquad b+b\implies \sqrt{8}+\sqrt{8}\)

OpenStudy (anonymous):

vertex is the center.. right

OpenStudy (jdoe0001):



OpenStudy (anonymous):

thanks for explaining.. but im a lost pup -_-

OpenStudy (jdoe0001):

yw